OpenStudy (anonymous):

(Can someone pleaseee pleasee please help me with this? it's almost the end of the semester and I really need help ) x y 2 5 6 15 18 45 a. yes; y= 0.4x b. yes; y= 10x c. yes; y= 2.5x no; y does not vary directly with x

OpenStudy (anonymous):

@Lyrae

OpenStudy (anonymous):

@mathmath333 help pleaseee

OpenStudy (anonymous):

@MakaylaTracy

OpenStudy (anonymous):

someoneeee please :(

OpenStudy (anonymous):

So basically, all you have to do is plug x in to see which one is correct. Let's try A. y = 0.4x So, we'd plug in x and y. 5 = 0.4(2) 5 = .0.8 No. So we can cross A out because it's not true. Let's try B. y = 10x 5 = 10(2) 5 = 20 No. So we can cross B out. Let's try C. y = 2.5x 5 = 2.5(2) 5 = 5 Yes. 15 = 2.5(6) 15 = 15 Yes. 45 = 2.5(18) 45 = 45 Yes. Your answer is C.

OpenStudy (anonymous):

oooo okiiii :o

OpenStudy (anonymous):

thankies. :) imma fan you <3 and is it okay if I ask a few more questions?

OpenStudy (anonymous):

@BunnyBree

OpenStudy (anonymous):

Sure. :)